# RuleRAG: Rule-Guided Retrieval-Augmented Generation with Language Models for Question Answering

### v2

@@ Abstract
Retrieval-augmented generation (RAG) footnotetext: * Corresponding Authors has shown promising potential in knowledge intensive question answering (QA). However, existing approaches only consider the query itself, neither specifying the retrieval preferences for the retrievers nor informing the generators of how to refer to the retrieved documents for the answers, which poses a significant challenge to the QA performance. To address these issues, we propose Rule-guided Retrieval-Augmented Generation with LMs, which explicitly introduces rules for in-context learning (RuleRAG-ICL) to guide retrievers to recall related documents in the directions of rules and uniformly guide generators to reason attributed by the same rules. The combination of queries and rules can be used as fine-tuning data to update retrievers and generators, achieving better rule-based instruction-following ability (RuleRAG-FT). Moreover, most existing RAG datasets were constructed without considering rules and Knowledge Graphs (KGs) are recognized as providing high-quality rules. Therefore, we construct five rule-aware RAG benchmarks for QA, RuleQA, based on KGs to stress the significance of retrieval and reasoning with rules. Experiments on RuleQA demonstrate RuleRAG-ICL improves the retrieval quality of +89.2% in Recall@10 and answer accuracy of +103.1% in Exact Match, and RuleRAG-FT yields more enhancement. In addition, experiments on four existing RAG datasets show RuleRAG is also effective by offering rules in RuleQA to them, further proving the generalization of rule guidance in RuleRAG. Code and RuleQA are at https://anonymous.4open.science/r/RuleRAG.
@@
Zhongwu Chen1, Chengjin Xu2∗, Dingmin Wang3∗, Zhen Huang1∗, Yong Dou1, Jian Guo2 1National Key Laboratory of Parallel and Distributed Computing, 1College of Computer Science and Technology, National University of Defense Technology 2International Digital Economy Academy (IDEA) 3Department of Computer Science, University of Oxford {chenzhongwu20, huangzhen, yongdou}@nudt.edu.cn, {xuchengjin, guojian}@idea.edu.cn, {dingmin.wang}@cs.ox.ac.uk
@@ 1 Introduction
However, although rules are common and valuable for the QA task, most existing RAG datasets were constructed without considering rules. KGs are widely known to provide question-answer pairs and high-coverage rules, so we newly construct five rule-aware QA benchmarks (RuleQA) from five KGs to offer rich data. The strong performance of KG rule mining algorithms also ensures the acquisition of rules with high confidence. In addition, the answers in RuleQA require reasoning based on rules rather than directly repeating retrieved facts like existing RAG datasets, so RuleQA is knowledge-intensive and more challenging. Experiments on RuleQA show that, under several retrieval and generation configurations, RuleRAG-ICL offers considerable performance gains with the individual guidance of rules by in-context learning and RuleRAG-FT achieves further improvements by fine-tuning retrievers and generators. RuleRAG-FT can also be extrapolated to unseen rules without retraining, confirming the superiority of rules to retrieve and reason in RAG. Moreover, we conduct experiments on four existing RAG datasets by adaptively equipping our mined rules for them. We find both RuleRAG-ICL and RuleRAG-FT can improve their performance, implying that apart from our constructed benchmarks RuleQA, it is also beneficial to introduce rules to existing RAG datasets.
@@ 4.1 Benchmarks and Setup of RuleRAG
Table 2: Performance comparison of RuleRAG-ICL and RuleRAG-FT with their variants and baselines. RG-DPR and RG-LLAMA2_7B represent rule-guided DPR and rule-guided LLAMA2_7B in RuleRAG-ICL. RGFT represents rule-guided fine-tuning in RuleRAG-FT. SSFT represents standard supervised fine-tuning. Standard Prompting does not have a retrieval stage, VE and CoK involve multiple search objects, which change several times, so there is no R@10. The best performance of RuleRAG-ICL and RuleRAG-FT are in bold.
@@ 5.1 Main Results
Table 2 shows the overall experimental results in the five rule-aware QA benchmarks detailedly and provides a comprehensive comparison between our proposed RuleRAG-ICL, RuleRAG-FT and all the baselines, under the concrete instantiation of DPR (Karpukhin et al., 2020) and LLAMA2_7B (Touvron et al., 2023) as retrievers and generators. As a baseline without retrieval, LLAMA2_7B using standard prompting can only refer to the knowledge it acquired during pre-training. Unsurprisingly, we notice that Standard Prompting (LLAMA2_7B) yields the worst relative and absolute results in all the five benchmarks, revealing that parametric knowledge in LLMs makes it hard to answer our factual queries. Furthermore, the results of Standard Prompting avoid the concern that the performance improvement of subsequent experiments comes from intrinsic knowledge in LLMs. This also gives a side note to the challenges of our constructed five benchmarks and motivates the introduction of rules.
@@
For RuleRAG-ICL (RG-DPR + LLAMA2_7B), introducing rules in the retrieval stage alone enhances DPR recall performance and improves the answer accuracy of the original LLAMA2_7B. RuleRAG-ICL (RG-DPR + RG-LLAMA2_7B) consistently surpasses Standard RAG across various metrics (+9.3 in R@10, +5.9 in EM and +3.2 in T-F1 on average absolute performance over all five benchmarks), achieving the improved performance. This confirms the sub-optimal ability of the current RAG and the effectiveness of our proposed dual rule-guided retriever and generator. For RuleRAG-FT, our proposed RGFT can amazingly improve performance by a significant margin (+45.7 in R@10, +24.2 in EM and +15.3 in T-F1 compared to the best performance of RuleRAG-ICL).
@@ 5.3 Further Analysis on Generators
In Figure 4, the x-axis is the ratio of the amount of fine-tuned data to the total amount of data in RGFT. The y-axis is the ratio of EM performance to the optimal one under DPR and LLAMA2_7B, with closer to 100% indicating stronger performance. The overall performance trend of the five benchmarks is that the larger the amount of fine-tuning data, the better the results. Yet, since the different properties of the rules in different benchmarks lead to different degrees of difficulty in learning, the growth of model performance under different benchmarks exhibits various characteristics.
Specifically, RuleQA-F has the most intuitive growth curve: It starts from low performance and slowly grows to the optimal performance, reflecting the increasing mastery of rules during the RGFT process of RuleRAG-FT. RuleQA-W yields relatively superior performance after just one-eighth of the total amount of the fine-tuning data and further improves subsequently. RuleQA-N achieves the best EM performance after three-eighths of the fine-tuning data and maintains flat. In contrast, the performance in RuleQA-Y fluctuates modestly at a very low level throughout the first half of the RGFT process (from one-eighth to five-eighths), and then sees a sudden surge in capability during the second half of the RGFT process (from six-eighths to the end). The EM performance in RuleQA-I fluctuates more dramatically: While realizing very large EM performance gains (ranking second in all the benchmarks), it undergoes several upward and downward drops before levelling off at the optimal performance. This suggests that RuleQA-I is the most challenging among our constructed five benchmarks. Moreover, from Table 2,  4 and  5, we find RuleRAG has the worst absolute performance in RuleQA-I compared to the other four benchmarks under the same LLMs, which also illustrates the challenge of our constructed RuleQA-I.
@@ 6 Rule Generalization
CoK with Rules 40.0 16.2 38.6 35.4
Table 3: The results of RuleRAG and CoK on four common datasets adaptively equipped with rules in RuleQA.
Generalization on More Datasets
Even though these datasets were constructed without adapting rules, RuleRAG still achieves consistent performance gains with the help of the rules in our constructed RuleQA. Specifically, following the framework of RuleRAG, existing datasets can be adaptively equipped with rules in RuleQA by calculating the relevance between candidate rules and queries. If some rules are highly relevant, they are introduced, otherwise no rules are introduced. Table 3 also compares the performance changes of CoK without and with our introduced rules, indicating that the guidance of rules is beneficial to the retrieval and generation accuracy of existing models. These results further confirm the effectiveness of our proposed rule-guided retrieval and generation in RAG for more comprehensive QA applications.

### v3

@@ Abstract
Retrieval-augmented generation (RAG) has shown promising potential in knowledge intensive question answering (QA). However, existing approaches only consider the query itself, neither specifying the retrieval preferences for the retrievers nor informing the generators of how to refer to the retrieved documents for the answers, which poses a significant challenge to the QA performance. To address these issues, we propose Rule-guided Retrieval-Augmented Generation with LMs, which explicitly introduces rules for in-context learning (RuleRAG-ICL) to guide retrievers to recall related documents in the directions of rules and uniformly guide generators to reason attributed by the same rules. The combination of queries and rules can be used as fine-tuning data to update retrievers and generators, achieving better rule-based instruction-following ability (RuleRAG-FT). Moreover, most existing RAG datasets were constructed without considering rules and Knowledge Graphs (KGs) are recognized as providing high-quality rules. Therefore, we construct five rule-aware RAG benchmarks for QA, RuleQA, based on KGs to stress the significance of retrieval and reasoning with rules. Experiments on RuleQA demonstrate RuleRAG-ICL improves the retrieval quality of +89.2% in Recall@10 and answer accuracy of +103.1% in Exact Match, and RuleRAG-FT yields more enhancement. In addition, experiments on four existing RAG datasets show RuleRAG is also effective by offering rules in RuleQA to them, further proving the generalization of rule guidance in RuleRAG. Code and RuleQA are at https://anonymous.4open.science/r/RuleRAG.
@@
Zhongwu Chen1, Chengjin Xu2∗, Dingmin Wang3∗, Zhen Huang1∗, Yong Dou1, Xuhui Jiang2, Jian Guo2 1National Key Laboratory of Parallel and Distributed Computing, 1College of Computer Science and Technology, National University of Defense Technology 2IDEA Research, International Digital Economy Academy 3Department of Computer Science, University of Oxford {chenzhongwu20, huangzhen, yongdou}@nudt.edu.cn, {xuchengjin, jiangxuhui, guojian}@idea.edu.cn, {dingmin.wang}@cs.ox.ac.uk
@@ 1 Introduction
However, although rules are common and valuable for the QA task, most existing RAG datasets were constructed without considering rules. KGs are widely known to provide question-answer pairs and high-coverage rules, so we newly construct five rule-aware QA benchmarks (RuleQA) from five KGs to offer rich data. The strong performance of KG rule mining algorithms also ensures the acquisition of rules with high confidence. In addition, the answers in RuleQA require reasoning based on rules rather than directly repeating retrieved facts like existing RAG datasets, so RuleQA is knowledge-intensive and more challenging. Experiments on RuleQA show that, under several retrieval and generation configurations, RuleRAG-ICL offers considerable performance gains with the individual guidance of rules by in-context learning and RuleRAG-FT achieves further improvements by fine-tuning retrievers and generators. RuleRAG-FT can be extrapolated to unseen rules without retraining, confirming the superiority of rules to retrieve and reason in RAG. We also introduce RuleRAG into an advanced model CoK to emphasize the paradigm of Rule-guided RAG is suitable for different RAG methods. Moreover, we conduct experiments on four existing RAG datasets. As a result, RuleRAG is beneficial for our constructed RuleQA and for introducing rules to existing RAG datasets. RuleRAG-CoK shows the attribution of the advanced RAG-based variant of RuleRAG under RuleQA and existing RAG methods.
@@ 4.1 Benchmarks and Setup of RuleRAG
RuleRAG-CoK RG-DPR RG-LLAMA2_7B - 5.1 17.9 - 2.6 14.5 - 8.4 32.2 - 11.8 26.1 - 9.2 25.7
Table 2: Performance comparison of RuleRAG-ICL, RuleRAG-FT and the variant of RuleRAG, RuleRAG-CoK. RG-DPR and RG-LLAMA2_7B represent rule-guided DPR and rule-guided LLAMA2_7B in RuleRAG-ICL. RGFT represents rule-guided fine-tuning in RuleRAG-FT. SSFT represents standard supervised fine-tuning. Standard Prompting does not have a retrieval stage, VE and CoK involve multiple search objects, which change several times, so there is no R@10. The best performance of RuleRAG-ICL and RuleRAG-FT are in bold.
@@ 5.1 Main Results
Table 2 shows the overall experimental results in the five rule-aware QA benchmarks detailedly and provides a comprehensive comparison between our proposed RuleRAG-ICL, RuleRAG-FT, the variant of RuleRAG, RuleRAG-CoK, and all the baselines, under the instantiation of DPR (Karpukhin et al., 2020) and LLAMA2_7B (Touvron et al., 2023) as retrievers and generators. As a baseline without retrieval, LLAMA2_7B using standard prompting can only refer to the knowledge it acquired during pre-training. Unsurprisingly, we notice that Standard Prompting (LLAMA2_7B) yields the worst relative and absolute results in all the five benchmarks, revealing that parametric knowledge in LLMs makes it hard to answer our factual queries. Furthermore, the results of Standard Prompting avoid the concern that the performance improvement of subsequent experiments comes from intrinsic knowledge in LLMs. This also gives a side note to the challenges of our constructed five benchmarks and motivates the introduction of rules.
@@
For RuleRAG-ICL (RG-DPR + LLAMA2_7B), introducing rules in the retrieval stage alone enhances DPR recall performance and improves the answer accuracy of the original LLAMA2_7B. RuleRAG-ICL (RG-DPR + RG-LLAMA2_7B) consistently surpasses Standard RAG across various metrics (+9.3 in R@10, +5.9 in EM and +3.2 in T-F1 on average absolute performance over all five benchmarks), achieving the improved performance. This confirms the sub-optimal ability of the current RAG and the effectiveness of our proposed dual rule-guided retriever and generator. For RuleRAG-FT, our proposed RGFT can amazingly improve performance by a significant margin (+45.7 in R@10, +24.2 in EM and +15.3 in T-F1 compared to the best performance of RuleRAG-ICL). In addition to Standard RAG-based RuleRAG-ICL and RuleRAG-FT, RuleRAG can also be applied to many advanced RAG-based models. As a variant of RuleRAG, RuleRAG-CoK introduces the idea of rule-guided RAG into CoK. The performance improvement achieved is attributed to our proposal.
@@ 5.3 Further Analysis on Generators
In Figure 4, the x-axis is the ratio of fine-tuned data to the total amount of data in RGFT. The y-axis is the ratio of EM performance to the optimal one under DPR and LLAMA2_7B, with closer to 100% indicating stronger performance. Since the different properties of the rules in different benchmarks lead to different degrees of difficulty in learning, the growth of model performance under different benchmarks exhibits various characteristics.
The performance in RuleQA-Y fluctuates modestly at a very low level throughout the first half of the RGFT process, and then sees a sudden surge in capability during the second half of the RGFT process. It is worth noting that the EM performance in RuleQA-I fluctuates more dramatically: While realizing substantial EM performance gains (ranking second in all the benchmarks), it undergoes several upward and downward drops before levelling off at the optimal performance. This suggests that RuleQA-I is the most challenging among our constructed five benchmarks. Moreover, from Table 2,  4 and  5, we find RuleRAG has the worst absolute performance in RuleQA-I compared to the other four benchmarks under the same LLMs, which also illustrates the challenge of the rules in RuleQA-I.
@@ 6 Rule Generalization
Generalization on More Datasets
Even though these datasets were constructed without adapting rules, RuleRAG still achieves consistent performance gains with the help of the rules in our constructed RuleQA. Specifically, following the framework of RuleRAG, existing datasets can be adaptively equipped with rules in RuleQA by calculating the relevance between candidate rules and queries. If some rules are highly relevant, they are introduced, otherwise no rules are introduced. Table 3 also compares the performance changes of an advanced RAG model CoK without and with our proposed RuleRAG, indicating that when CoK replaces Standard RAG as the base method, the variant of RuleRAG, RuleRAG-CoK, still succeeds in introducing the guidance of rules. These results further confirm the effectiveness of our proposed rule-guided retrieval and generation in RAG for more comprehensive QA models and applications.
@@
RuleRAG-CoK 40.0 16.2 38.6 35.4
Table 3: The results of Standard RAG and CoK on four RAG datasets before and after equipping RuleRAG.